The Peng-Robinson EOS is a function which describes the absolute pressure of a gas as a

function of the absolute temperature T, the molar volume, V, and several constants:

P = RT/(V-b) - a / [V(V+b) + b(V-b)]

A set of unit-consistent constants for this equation are: R = 83.14 bar-cm^3/(mol*K) (i.e. the Gas Constant), a = 2.417x10^7 bar*cm^6/mol^2 and b = 90.18 cm^3/mol. Note: these two values are fitted to a particular fluid, in this case, n-pentane.

Plot f(V) = P - {RT/(V-b) - a / [V(V+b) + b(V-b)]} for P = 1 bar and T = 100 °C using Excel, choosing the lower and upper bounds of V = [130, 500]. Choose a logical interval for V in

between these two limits (ex. 5 or 10). Be sure to convert T to an absolute scale.
Answer #1

Absolute Temperature in Kelvin = Temperature in C + 273.15

The given function is directly plotted using excel.

24170000 bar*cm6/mol^2 90.18 cm^3/mol 1 Bar 373.15 K 83.14 (Absolute Scale) Volume f(V) 130 -27.811876 140 36.5452385 150 65

The formula version is :

-2.417 * 1017 =90.18 bar*cm^6/mol^2 cm^3/mol Bar (Absolute Scale) = (100+273.15) 83.14 Volume 130 =A10 + 10 =A11 + 10 =A12 +

f(V) vs Volume f(V) 100 200 300 400 500 600 Volume (V)